OpenStudy (anonymous):

When tossing 2 coins, P(at least 1 tail) = 1 - P(not all tails).

OpenStudy (kinggeorge):

Is this a true or false? What's the question?

OpenStudy (anonymous):

yeah its true or false

OpenStudy (kinggeorge):

Do you have a guess before I explain it?

OpenStudy (anonymous):

false

OpenStudy (kinggeorge):

That's correct. I like to think about it rather explicitly. P(not all tails)=P(0 tails or 1 tail). Thus, 1-P(0 tails or 1 tail)=P(2 tails) since you're tossing two coins. This is obviously not the same as P(1 tail).

OpenStudy (anonymous):

Thank you :)

OpenStudy (kinggeorge):

You're welcome.